# Do AI Models Perform Human-like Abstract Reasoning Across Modalities?

### v3

@@ 3.1 Output Grid Accuracy
Table 1: Reasoning models: Output-grid accuracy (pass@1) for Concept-ARC across models and experimental settings. Accuracy is shown in %. Each cell shows textual / visual accuracy. For o3 and o4-mini, we use the “low” and “medium” effort settings in the OpenAI API. For Claude and Gemini, we use a 16K reasoning token budget to approximate o3’s medium effort setting. Temperature is set to 1 for all models. Bold numbers correspond to the highest visual and textual scores in each column.
Reasoning model low effort medium effort low effort + tools medium effort + tools
Textual / Visual Textual / Visual Textual / Visual Textual / Visual
o3 68.3 / 6.7 77.1 / 5.6 67.9 / 18.1 75.6 / 29.2
o4-mini 52.1 / 3.8 70.8 / 8.1 57.3 / 6.7 77.7 / 25.0
Claude Sonnet 4 N/A 60.2 / 5.2 N/A 55.0 / 6.9
Gemini 2.5 Pro N/A 66.0 / 4.2 N/A 60.4 / 5.8
@@
Table 1 and Table 4 (in the Appendix) give the pass@1 output-grid accuracies of the reasoning models and non-reasoning models we evaluated in both textual and visual modalities. [^7] Following the ARC-Prize evaluation ARC-Prize (2024), we counted an output grid as correct only if it perfectly matched the ground truth and was in the requested format. Upon examination of the output grids generated by different models, we found that in some cases models generated these grids in other formats, such as rows enclosed by brackets, and values separated by commas. If we accept all different grid formats, we find that accuracies in the textual modality in Table 1 improve in some cases, whereas accuracies in the visual modality do not change. We give a more complete description in Appendix I. In all cases, non-reasoning models attain much lower accuracy than reasoning models, so here we focus our analysis on the reasoning models. For all models, we see a dramatic performance gap between the textual and visual settings. Further, especially for o3 and o4-mini, and to a lesser extent for Claude and Gemini, we observe a jump in visual accuracy when Python tools are enabled. In contrast, allowing Python tools does not have a similar effect in the textual setting for three of the models, with o4-mini being the only exception. For o3 and o4-mini, increased reasoning effort is associated with increased accuracy in the textual modality, with or without tools; in the visual modality, we observe that the models primarily use the increased reasoning budget to execute more Python code, which may explain the substantial improvement in medium effort + tools.
@@
Inspecting the failure cases of the visual setting more closely, we find that models struggle to recognize the correct grid size from the image inputs. When Python tools are enabled, the models use computer vision libraries to partially compensate for this difficulty. In both textual and visual modalities, the majority of incorrect output grids are due to a simple mismatch between the generated and ground-truth grids, but—particularly in the visual setting—there is a small share of invalid outputs, either due to uneven row lengths or non-integer tokens in the grid. Figure 6 gives an error-type distribution for o3.
@@
Using unpublished data from Moskvichev et al.’s 2023 study, we found that human-generated output grids achieved an overall pass@1 accuracy of 73% on the 480 ConceptARC tasks, lower than that of the top reasoning models in the textual modality. (We provide per-concept accuracies in Appendix F.)
3.2 Rule Evaluation
@@
Our team manually evaluated the rules generated by o3 in all settings and by Claude Sonnet 4 and Gemini 2.5 Pro in the medium-effort + tools setting, for both textual and visual modalities. We also evaluated the pass@1 rules generated by humans, using data from the study by Moskvichev et al. (2023). Each rule’s assessment was agreed upon by at least two members of our team. Due to low accuracy of other settings and resource limitations of our team, we did not evaluate rules from other models or experimental settings. However, the evaluations we performed provide substantial insight into the ability of different models (and the effects of different model settings), as well as human participants, to grasp intended abstractions of ConceptARC tasks.
@@
Figure 2 shows the results of our rule evaluations on o3, Claude, and Gemini, all using medium-effort + tools, in both textual and visual modalities, as well as evaluations for human-generated rules. Specifically, there are two bars for each model: Correct Grid and Incorrect Grid. The height of each bar corresponds to the percentage of the 480 tasks on which the model’s output grid was correct or incorrect. Within each bar, the green section corresponds to tasks for which the model’s generated rule was correct-intended; the yellow section corresponds to correct-unintended rules; and the red section corresponds to incorrect rules. The left section of the plot gives these results for the textual modality and the middle section for the visual modality. The rightmost section gives the output accuracy and rule evaluation results for human-generated rules. The gray areas in the human-result bars correspond to human-generated solutions for which we were unable to classify the rule, either because no rule was given by the participant, no rule was collected by the experimenters (this was the case for all of the Incorrect tasks), or the rule given was too unclear to confidently evaluate.
@@
Notably, while o3 in the textual setting rivals humans in output grid accuracy, about 28% of its correct outputs are based on correct-unintended or incorrect rules—indicating reasoning based on superficial patterns rather than intended abstract concepts. We found several types of unintended rules used by models, including rules that described complicated (and spurious) patterns in the demonstrations, rules that focused on irrelevant features such as (in the textual setting) the specific numbers encoding grid colors, and rules that came close to capturing intended abstractions but included irrelevant spurious associations. Examples of such rules are given in Figure 4. In comparison, we found that only 8% of human’s correct outputs were similarly based on correct-unintended or incorrect rules. While our analysis for human-generated rules is limited due to missing rule data (about 20% of rules for correct outputs were not classifiable), this difference is suggestive and should be clarified in future research. Comparing AI models with one another, in cases where Claude and Gemini were accurate on output, both have a smaller fraction of correct-unintended rules than o3, but both are lower than o3 in output accuracy.
@@
Also notable is the percentage of incorrect output grids that are based on correct-intended rules. In these cases, the models recognized the intended abstract rule describing the grid transformation, but were unable to apply it correctly to the test grid. In the textual setting, this seems to be most common in Claude, and less so in Gemini and o3. In the visual setting, however, o3 produced correct-intended rules in about 27% of cases in which its output grid was incorrect; Claude and Gemini did so less frequently, but both still at substantial rates. In summary, looking only at a model’s output accuracy in the textual setting—as was done in (Chollet, 2024)—might overestimate the model’s ability for abstract reasoning, but in the visual domain, accuracy alone might underestimate its abstract reasoning abilities. This hints at a direction for improvement in the visual modality across models: models with the capacity to apply the determined rule correctly would be able to substantially improve their output accuracy. These insights illustrate the importance of going beyond simple accuracy in assessing the capabilities of AI models.
@@
While Figure 2 showed our rule evaluations for different models using medium reasoning effort + tools, Figure 3 shows the effects of varying reasoning effort and tool use for the o3 model, in both textual and visual modalities. There are a few important observations to make. First, in the visual setting, increasing reasoning effort from low to medium alone does not have any substantial effect on output accuracy or rule correctness, which aligns with prior work suggesting that test-time scaling does not have the dramatic effects in visual modalities that have been seen in text-only LLMs (Hao et al., 2025). However, enabling Python tool use does result in substantial improvement in output accuracy and rule correctness, especially at medium reasoning effort, likely because the model is able to use computer vision libraries. In contrast, in the textual setting, increasing reasoning effort has a larger positive effect on both output accuracy and rule correctness than enabling Python tool use.
@@
Figure 2: Results of rule evaluations. For each model in each modality (as well as humans), two bars are given, representing the percentage of correct and incorrect grid outputs over the 480 ConceptARC tasks. Each bar shows the fraction of tasks for which the rule is correct-intended, correct-unintended, and incorrect. The gray areas in the human-result bars represent rules that we could not classify—see subsection 3.2 for details. The actual percentages corresponding to regions on the bars are given in Appendix D.
@@
Figure 3: Results of rule evaluations for o3 across all settings. As in Figure 2, two bars showing the percentage of correct and incorrect output grids are included for each setting, with each bar showing the fraction of tasks for which the generated rule is correct-intended, correct-unintended, and incorrect. The actual percentages corresponding to regions on the bars are given in Appendix D.
Figure 4: Examples of correct-unintended rules. Top: o3, using medium effort and tools, performs shallow inference for a task from the Horizontal vs. Vertical concept group. The model does not recognize the relation between the orientation of the colored shape components and the blue row, but rather focuses on whether a blue (“8”) pixel appears in the grid. In this case, the correct-intended rule works for the given test case, but does not work for other test variants. Middle: o3, using medium effort and tools, on a task from the Complete Shape concept group. The model does not recognize the relation between the colored output shape and the gray prototype and instead overfits to the training examples, producing a correct-intended rule based on shallow features. Bottom: Claude Sonnet 4 uses a density heuristic to approximate the most overlapped figure on a task from the Top vs. bottom 3D group. While this works for some of the test examples, it does not capture the notion of the bottommost shape in a 3D stack, and there are several possible scenarios for which this approach fails.
@@ 4 Discussion
Given the results described above, we can now provide preliminary answers to the questions we listed at the beginning of this paper. (1) How does the accuracy obtained by AI models compare with that of humans? Table 1 shows that for textual inputs, o3, with medium reasoning effort, matches or surpasses human accuracy on ConceptARC tasks, with Claude and Gemini obtaining lower accuracy, and o4-mini surpassing humans only when Python tools are enabled. This aligns with results reported in (Chollet et al., 2025, ARC-Prize, 2025).888(Kamradt, 2025) noted the large discrepancy between the accuracy of o3-preview (the pre-release version) and the released version of o3 on ARC-AGI-1. The reasons for this discrepancy are not known. However, using the visual modality, the models’ performance still lags significantly behind human accuracy, even when models are given access to Python tools.
@@
(2) To what extent do the rules generated by AI models capture the abstractions that were intended by ConceptARC’s creators, versus more superficial shortcuts? Figure 2 shows that for textual inputs and medium reasoning effort with Python tools, about 57% of o3’s generated rules (regardless of output accuracy) were correct and intended; that is, they captured the intended abstractions of the tasks. However, about 28% of o3’s generated rules were correct but unintended, meaning they were correct with respect to the given demonstrations, and frequently generated correct output grids, but did not capture the intended abstractions. ConceptARC, like ARC, is built on “core knowledge” priors, including “objectness” Chollet (2019), but we found that, for example, o3’s rules often focused on colors and individual pixels rather than objects. Moreover, using integers to encode colors enabled unintended shortcuts such as relying on numerical values (e.g., the value for green, 3, is greater than the value for red, 2) which were not available in visual modalities. Both Claude and Gemini’s shares of correct-unintended rules (14% and 17% respectively) were lower than o3’s, but more than twice the percentage of correct-unintended rules produced by humans (3%). Thus AI models seem more likely to miss intended abstractions and to solve tasks using more superficial features than humans.
@@
(3) Regarding the effects of textual vs. visual modalities, Table 1 and Figure 2 show that both output-grid and rule correctness drop dramatically in the visual mode. In addition, we observe that in this mode all three models are considerably better at forming correct-intended rules than generating correct output grids. As for the effects of reasoning effort and Python tools, Table 1 and Figure 3 show that the former is more helpful for textual inputs and the latter is more helpful for visual inputs, especially at higher reasoning effort. These results point to possible directions for strengthening visual reasoning models, especially in more abstract domains.
@@
In short, our results show that models still lag humans in abstract reasoning, and that using accuracy alone to evaluate abstract reasoning on ARC-like tasks may overestimate abstract-reasoning capabilities in textual modalities and underestimate it in visual modalities. In evaluating capabilities such as abstract reasoning in AI systems, our results highlight the importance of going beyond simple accuracy, namely assessing both robustness and the extent to which a system uses generalizable mechanisms rather than more superficial shortcuts (Frank, 2023, Ivanova, 2025, Rane et al., 2025). More generally, developing AI models that are better at grasping the abstractions understood by humans will be essential for these systems to generalize in human-like ways, and for these systems to be able to explain their reasoning in ways understandable to humans—both key abilities for successful human-AI interaction.

### v4

@@ 3 Results
Table 1: Reasoning models: Output-grid accuracy (pass@1) for ConceptARC across models and experimental settings. Accuracy is shown in %. Textual (t) and Visual (v) results are reported on separate rows. For o3 and o4-mini, we use the “low” and “medium” effort settings in the OpenAI API. For Claude we use the model Sonnet-4 and for Gemini, we use 2.5 Pro. For both models, we use a 16K reasoning token budget to approximate o3’s medium effort setting. We denote enabled tool usage with +T. Temperature is set to 1 for all models. Bold numbers correspond to the highest visual and textual scores in each column.
Model low medium low +T medium +T
o3 t 68.3 77.1 67.9 75.6
v 6.7 5.6 18.1 29.2
o4-mini t 52.1 70.8 57.3 77.7
v 3.8 8.1 6.7 25.0
Claude t N/A 60.2 N/A 55.0
v N/A 5.2 N/A 6.9
Gemini t N/A 66.0 N/A 60.4
v N/A 4.2 N/A 5.8
Output Grid Accuracy
@@
Table 1 and Table 3 (Appendix E) give the pass@1 output-grid accuracies of the reasoning models and non-reasoning models we evaluated in both textual and visual modalities. In all cases, non-reasoning models attain much lower accuracy than reasoning models, so here we focus our analysis on the reasoning models. For all models, we see a dramatic performance gap between the textual and visual settings. Further, especially for o3 and o4-mini, and to a lesser extent for Claude and Gemini, we observe an increase in visual accuracy when Python tools are enabled. In contrast, allowing Python tools does not have a similar effect in the textual setting for three of the models, with o4-mini being the only exception. For o3 and o4-mini, increased reasoning effort is associated with increased accuracy in the textual modality, with or without tools; in the visual modality, we observe that the models primarily use the increased reasoning budget to execute more Python code, which may explain the substantial improvement in medium effort + tools.
@@
Inspecting the failure cases of the visual setting more closely, we find that models struggle to recognize the correct grid size from the image inputs. When Python tools are enabled, the models use computer vision libraries to partially compensate for this difficulty. In both textual and visual modalities, the majority of incorrect output grids are due to a mismatch between the generated and ground-truth grids; however, in the visual setting in particular, there is also a small share of invalid outputs, either due to uneven row lengths or non-integer tokens in the grid. Figure 4 in Appendix G gives an error-type distribution for o3.
@@
Using unpublished data from Moskvichev et al.’s (2023) study, we found that human-generated output grids achieved an overall pass@1 accuracy of 73% on the 480 ConceptARC tasks, lower than that of o3 in the textual modality. (We provide per-concept accuracies in Appendix H.)
Rule Evaluation
@@
Our team manually evaluated the rules generated by o3 in all settings and by Claude Sonnet 4 and Gemini 2.5 Pro in the medium-effort + tools setting, for both textual and visual modalities. We also evaluated the pass@1 rules generated by humans, using data from the study by Moskvichev et al. (2023). For each rule (human or machine-generated), an initial classification was assigned by one member of our team, and reviewed by a second member. In cases where there was disagreement or uncertainty about a rule’s classification, our team discussed the rule together until we came to a consensus. Due to low accuracy in other settings and resource limitations of our team, we did not evaluate rules from other models or experimental settings. Nevertheless, the selected evaluations provide substantive insight into the extent to which different models and human participants utilize the intended abstractions of ConceptARC tasks.
@@
Figure 2 shows the results of our rule evaluations on o3, Claude, and Gemini, all using medium-effort + tools, in both textual and visual modalities, as well as evaluations for human-generated rules. In each setting there are two bars for each model: Correct Grid and Incorrect Grid. The height of each bar corresponds to the percentage of the 480 tasks on which the model’s output grid was correct or incorrect. Within each bar, the green, yellow, and red sections correspond respectively to tasks for which the model’s generated rule was correct-intended, correct-unintended rules, and incorrect; gray sections correspond to unclear or nonresponsive rules. Results for the textual and visual modalities are given on the left and the middle parts respectively. The rightmost part gives the output accuracy and rule evaluation results for human-generated rules. The gray areas in the bars correspond to solutions for which we were unable to classify the rule, either because no rule was given by the participant or model, no rule was collected by the experimenters (this was the case for all of the human tasks with incorrect outputs), or the rule given was too unclear to confidently evaluate.
@@
Notably, while o3 in the textual setting rivals humans in output grid accuracy, around 27% of its correct outputs are based on correct-unintended or incorrect rules, indicating reasoning based on superficial patterns rather than intended abstract concepts. We found several types of unintended rules used by models, including rules that described complicated (and spurious) patterns in the demonstrations, rules that focused on irrelevant features such as numerical relationships among the specific numbers encoding grid colors, and rules that came close to capturing intended abstractions but included irrelevant spurious associations. In comparison, we found that only about 8% of humans’ correct outputs were based on correct-unintended or incorrect rules. Although our analysis for human-generated rules is limited due to missing rule data (roughly 19% of the rules associated with correct outputs were not classifiable, and no rules were collected for incorrect outputs), this difference is suggestive and should be clarified in future human studies. Comparing AI models with one another, in cases where Claude and Gemini generated accurate output grids, both have a smaller fraction of correct-unintended rules than o3, but both are lower than o3 in output accuracy.
@@
Also notable is the percentage of incorrect output grids that are based on correct-intended rules. In these cases, the models recognized the intended abstract rule describing the grid transformation, but were unable to apply it correctly to the test grid. In the textual setting, this seems to be most common in Claude, and less so in Gemini and o3. In the visual setting, however, o3 produced correct-intended rules in around 28% of cases in which its output grid was incorrect; Claude and Gemini did so less frequently, but both still at substantial rates. In summary, looking only at a model’s output accuracy in the textual setting (as was done by Chollet et al. (2024)) might overestimate the model’s ability for abstract reasoning, but in the visual domain, accuracy alone might underestimate its abstract reasoning abilities. This hints at a direction for improvement in the visual modality across models: models with the capacity to apply the determined rule correctly would be able to substantially improve their output accuracy. These insights illustrate the importance of going beyond simple accuracy in assessing the capabilities of AI models.
@@
Whereas Figure 2 showed our rule evaluations for different models using medium reasoning effort + tools, Figure 3 shows the effects of varying reasoning effort and tool use for the o3 model, in both textual and visual modalities. There are a few important observations to make. First, in the visual setting, increasing reasoning effort from low to medium alone does not have any substantial effect on output accuracy or rule correctness, which aligns with prior work suggesting that test-time scaling does not have the dramatic effects in visual modalities that have been seen in text-only models (Hao et al., 2025). However, enabling Python tool use does result in substantial improvement in output accuracy and rule correctness, especially at medium reasoning effort, likely because the model is able to use computer vision libraries. In contrast, in the textual setting, increasing reasoning effort has a larger positive effect on both output accuracy and rule correctness than enabling Python tool use.
Table 2: Percentage of tasks in which models exhibited either “rule–grid alignment” (RGA)—the generated rule accurately described the generated output grid—or, in visual settings, a “visual error” (VE). The medium-effort + tools setting was used for all models. Tasks with missing or invalidly formatted output grids and/or non-responsive or unclear rules are excluded from these percentages.
Model Textual Visual
(Medium + Tools)
RGA RGA VE
o3 98.1 97.4 49.1
Claude 92.4 96.6 59.7
Gemini 94.5 93.6 77.3
@@
Figure 2: Results of rule evaluations. For each model, as well as for humans, two bars are shown, representing the percentages of correct and incorrect grid outputs across the 480 ConceptARC tasks. Model evaluations are reported for both textual and visual settings. Each bar shows the fraction of tasks for which the rule is correct-intended, correct-unintended, and incorrect. The gray regions represent rules that could not be classified, further described in the section on Rule Evaluation. The exact percentages are given in Appendix I.
@@
Figure 3: Results of rule evaluations for o3 across all settings. As in Figure 2, two bars showing the percentage of correct and incorrect output grids are included for each setting, with each bar showing the fraction of tasks for which the generated rule is correct-intended, correct-unintended, and incorrect. Gray regions represent rules that could not be classified; see the section on Rule Evaluation for details. The exact percentages are given in Appendix I.
Rule-Grid Alignment
To investigate how accurately the natural-language rules reflect the models’ underlying reasoning, our team manually evaluated the solutions generated by o3, Claude Sonnet 4, and Gemini 2.5 Pro in the medium-effort + tools setting for two additional features: rule-grid alignment and visual errors. Rule-grid alignment describes cases where the model’s stated rule accurately captures the transformation demonstrated in its output grid. Visual errors refer to cases in visual settings where a grid’s incorrectness or inconsistency with its corresponding rule appears to stem, at least in part, from failures in the model’s perceptual apparatus rather than from a mismatch between its generated rule and underlying reasoning. Such errors commonly involved failures to recognize the exact grid dimensions, slight inaccuracies in object placement, or incorrect mappings between colors and their numerical encodings.
For these evaluations, one team member provided an initial judgment for each task, after which any ambiguous cases were discussed as a group until consensus was reached. The results of these evaluations are presented in Table 2. We find that the natural-language rules align with their corresponding grids in the vast majority of cases; across all evaluated models and settings, agreement exceeded 90% of tasks. This supports the view that the proposed rules generally reflect the reasoning used to produce the grid solutions. In addition, the relatively high rate of visual errors corroborates our earlier observations of failure modes in visual settings. Even the best-performing model and setting exhibited a visual error rate approaching 50%, despite maintaining a high degree of rule-grid alignment. This suggests that the low grid accuracies observed in visual settings may be driven in large part by perceptual limitations, rather than solely by differences in reasoning capacity between visual and textual modalities.
@@ 4 Discussion
We can now provide preliminary answers to the questions we listed at the beginning of this paper. (1) How does the accuracy obtained by AI models compare with that of humans? Table 1 shows that for textual inputs, o3 with medium reasoning effort surpasses human accuracy on ConceptARC tasks, with Claude and Gemini obtaining lower accuracy, and o4-mini surpassing humans only when Python tools are enabled. This aligns with results reported in Chollet et al. (2025) and ARC-Prize (2025). However, using the visual modality, the models’ performance still lags significantly behind human accuracy, even when models are given access to Python tools.
@@
(2) To what extent do the rules generated by AI models capture the abstractions that were intended by ConceptARC’s creators, versus more superficial shortcuts? Figure 2 shows that for textual inputs and medium reasoning effort with Python tools, about 57% of o3’s generated rules (regardless of output accuracy) were correct and intended; that is, they captured the intended abstractions of the tasks. However, about 29% of o3’s generated rules were correct but unintended, meaning they were correct with respect to the given demonstrations, and frequently generated correct output grids, but did not capture the intended abstractions. ConceptARC, like ARC, is built on “core knowledge” priors, including “objectness” (Chollet, 2019); however, we found that for the AI models we studied, rules often focused on colors, individual pixels, and other low-level features rather than objects. Moreover, using integers to encode colors enabled unintended shortcuts, such as relying on relationships between numerical values. Both Claude and Gemini’s shares of correct-unintended rules (approximately 15% and 22%, respectively) were lower than o3’s, but more than five times that of correct-unintended rules produced by humans (2.7%). Thus, AI models appear much more likely than humans to miss intended abstractions and instead solve tasks using more superficial features.
@@
(3) Regarding the effects of textual versus visual modalities, Table 1 and Figure 2 show that both output-grid and rule correctness drop dramatically in the visual mode. In addition, we observe that in the visual mode, all three models are considerably better at forming correct-intended rules than at generating correct output grids. Regarding the effects of reasoning effort and Python tools, Table 1 and Figure 3 show that increased reasoning effort is more helpful for textual inputs, whereas Python tools are more helpful for visual inputs, especially at higher reasoning effort. These results point to possible directions for strengthening visual reasoning models, especially in more abstract domains.
@@
In short, our results show that models still lag behind humans in the kinds of abstract reasoning capabilities that ARC and ConceptARC were designed to evaluate. Using accuracy alone to evaluate abstract reasoning on ARC-like tasks may overestimate abstract reasoning capabilities in textual modalities and underestimate them in visual modalities. Our results highlight the importance of going beyond simple accuracy, namely, assessing both robustness and the extent to which a system relies on generalizable mechanisms rather than superficial shortcuts (Frank, 2023; Ivanova, 2025; Rane et al., 2025). To target these abilities in a meaningful way, we encourage designers of benchmarks and evaluation methods to incorporate underlying abstractions and derived rules, alongside traditional output correctness, when developing new performance metrics. More generally, while in some cases it is desirable for AI systems to reason in non-human-like ways (e.g., in discovering novel patterns in scientific data, such as protein sequences (Jumper et al., 2021)), developing AI models that grasp the abstractions understood by humans will be essential for these systems to generalize and explain their reasoning in ways understandable to humans, which are key abilities for successful human-AI interaction. In future work, we will investigate whether human-like reasoning can be induced via process-based reward models or a more direct inclusion of human-generated reasoning traces. An interesting direction for future research will be to extend our studies to tasks that require more compositional reasoning, such as those in ARC-AGI-2 (Chollet et al., 2025).
5 Related Work
Many studies have proposed computational methods for solving ARC tasks, including program synthesis with and without pretrained LLMs (Banburski et al., 2020; Bober-Irizar and Banerjee, 2024; Wind, 2020), fine-tuning LLMs with augmented data and test-time training (Chollet et al., 2024), and using fine-tuned large reasoning models (Chollet, 2024). Almost all of these studies rely on textual task representations (Hu et al., 2025, is an exception), and all focus exclusively on output-grid accuracy; to our knowledge, none analyze the rules generated by models, as was done in our study. LeGris et al. (2025) collected rules generated by humans and trained a Naive Bayes classifier to predict which task in their study was described by a particular rule, finding that humans indeed used expected core knowledge priors (i.e., concepts related to objectness, basic geometry and topology, and numerosity) in their rules.
Besides ARC and ConceptARC, several benchmark datasets have been used to evaluate abstract and visual reasoning abilities in LLMs and large reasoning models. Those closest to ARC and ConceptARC include Bongard problems (Bongard, 1970), letter-string analogies (Hofstadter, 1985), Raven’s progressive matrices (RPMs) (Raven, 1938), and compositional visual reasoning (CVR) (Zerroug et al., 2022). Bongard problems are similar to ConceptARC tasks in that each problem tests understanding of a single core spatial or semantic concept, such as “large vs. small” or “inside vs. outside.” Bongard problems are inherently visual, and several studies have evaluated multimodal models on subsets or variations of the original problems, finding that, consistent with our results on ConceptARC, the poor performance of VLMs on these problems appears to arise primarily from difficulties with vision rather than with reasoning (Małkiński et al., 2024; Pawlonka et al., 2025; Wüst et al., 2024). Letter-string analogies were first proposed by Hofstadter (1985) as an idealized domain for analogy-making. Webb et al. (2023) found that GPT-3 reached human level accuracy on letter-string analogies, but other studies (testing both GPT-3 and GPT-4) found that LLMs were not robust to problem variations that did not affect humans’ performance (Hodel and West, 2023; Lewis and Mitchell, 2025). Raven’s progressive matrices (RPMs) have long been used as tests of human fluid intelligence. RAVEN, a dataset consisting of programmatically generated, simplified RPMs (Zhang et al., 2019), has been used to evaluate visual reasoning in VLMs (e.g., (Zhang et al., 2024; Zhu et al., 2025)), which seem to struggle more with perceptual understanding than with reasoning. Zerroug et al. (2022) proposed the CVR visual reasoning benchmark, which tests compositional reasoning abilities based on concepts similar to those used in ConceptARC. CVR tasks have been used to evaluate convolutional neural networks as well as vision transformers, none of which approach human-level performance.